# Walk Wisely on Graph: Knowledge Graph Reasoning with Dual Agents via Efficient Guidance-Exploration

### v1

@@ Conclusion
Cluster State Similarity
To assess the exploration efficiency of GIANT, we introduce cluster state similarity (CSS) as a metric, and in the main content we compare the performance of CURL and FULORA on only three tasks due to space constraints. Figure 8 also compare the model performance on remaining six tasks. For the same task, CURL experiences significant oscillations as path length increases, while FULORA remains stable. As in our previous analysis, CURL can only judge path quality by whether it has gone to the correct target cluster, while FULORA utilizes dynamic path feedback to promote the GIANT to converge to a high-quality path.
Figure 9: The difference value to ratios of the mean and variance of CSS between FULORA and CURL on different task-path_length.
Figure 9 further visualizes performances of FULORA and CURL on these tasks with varied path length. We utilize the ratio of the mean to the variance of CSS as an evaluation metric for the mean measures inference accuracy and the variance measures inference stability. In the vast majority of cases, FULORA outperforms CURL. In the case of long-distance reasoning for complex tasks like PersonBorninLocation-Path9 (PBL-9), AthletePlaysinLeague-Path7 (APL-9), AthletePlaysSport-Path9 (APS-9), and TeamPlaysinLeague (TPL), the performance gap between FULORA and CURL is significant.
Additional Ablation Studies
Here, we conduct a series of ablation experiments designed to answer the following three questions:
Q1: Can attention mechanism enhance DWARF’s reasoning ability?
Q2: Can dynamic path feedback accelerate DWARF’S learning speed via improving GIANT’S reasoning ability?
Q3: Can the efficient guidance-exploration approach make the guidance-exploration trade-off?
With a slight abuse of abbreviation, we use ATTN, DPF and GE to denote attention mechanism, dynamic path feedback and the efficient guide-exploration approach respectively. To highlight the effect of each component, we set FULORA as the baseline. In Figure 10, FULORA-X represents pulling module X out of the FULORA framework. We conduct experiments on three different scales of KG to make a fair comparison. It is worth noting that in order to describe DWARF reasoning accuracy in the whole training process, Entity State Similarity (ESS, be similar to CSS we mentioned above) is used as an evaluation index.
Attention Mechanism
@@
Dynamic Path Feedback
@@
Efficient Guidance-Exploration
The core of FULORA is the efficient guidance-exploration method, which contributes to make the guidance-exploration trade-off. In the previous HRL-based reasoning method like CURL, there is a high degree of coupling between the two agents. We treat ESS and CSS as two time series because they change with training progress. Next, we use Granger Causality test to examine the degree of coupling between DWARF and GIANT.
@@
The results of Granger Causality test for FULORA and FULORA-GE are shown in Table 6. Efficient guidance-exploration method reduces the coupling between GIANT and DWARF, thereby alleviating the affect of distribution deviation caused by cluster mapping. In particular, on the WN18RR and FB15K-237, FULORA performs better than the version without efficient guidance-exploration method due to the greatly reduced coupling between DWARF and GIANT. Certain tasks in WN18RR (sparse KG) require the agent to have excellent long-distance reasoning ability. However, due to the distribution deviation, even if the GIANT reaches the correct target cluster, the guidance provided by it may not be suitable for DWARF (for example, the path length is too long to reach the correct target entity). The high coupling causes GIANT also seriously affected by DWARF, resulting in poor model performance.
@@
Appendix C C. Time Series Analysis to Efficient Guidance-Exploration
Before Granger Causality test, we firstly use Augmented Dickey-Fuller (ADF) test to judge the stationarity of the series. The results as shown in Table 7. Next, we make D-order difference for the non-stationary time series, and then test the stationarity of the difference series. As illustrated in Figure 11, the time series is stationary after 1-order difference. Therefore, we demonstrate that these series are integrated of order one. To avoid the pseudo-regression phenomenon, we also need to do cointegration analysis of these series (Dougherty 2011). Here, we employ Johansen test based on maximum likelihood estimation for cointegration analysis.
Figure 11: Entity state similarity curves, cluster state similarity curves and their 1-order difference on WN18RR and FB15K-237 from FULORA and FULORA-GE.
In the Johansen cointegration test, the 2-cointegration term critical values of {10%, 5%, 1%} are {2.705, 3.841, 6.635} respectively. The critical value 3.841 corresponding to 5% was selected as the boundary for testing. Table 8 shows the results of Johansen cointegration test for FULORA and FULORA-GE. It shows that the seven non-stationary series avoid the pseudo-regression problem, thus the result of Granger causality test is reasonable.
@@
DWARF(-GE) $\leftrightarrow$ GIANT(-GE) 0.037 74.582 0.002 4.133
@@
Appendix D D. Proof of Dynamic Path Feedback
A major concern with dynamic path feedback is whether the optimal policy GIANT learns is consistent with the default rewards. Recall Eq.(12) where the next state $s^{c}_{t+1}$ is used, which means that it implies the action $a^{c}_{t}$. The next state $s^{c}_{t+1}$ is the output of the current state $s^{c}_{t}$ and the current action $a^{c}_{t}$ to the transition function. In fact, $J(\theta_{\pi^{c}})$ is a function of state $s^{c}_{t}$ and action $a^{c}_{t}$. The same goes for reward settings in the default environment. Therefore, we use the Q-function to simplify our notation,
@@
$=\mathbb{E}_{s_{t+1}^{c}\sim P(\cdot|s_{t}^{c},a_{t}^{c})}\left[r_{c}(s_{t}^{c})+\underset{a_{t+1}^{c}\in\mathcal{A}}{\text{max}}Q^{*}(s^{c}_{t+1},a^{c}_{t+1})\right],$
@@
$Q^{*}(s_{t}^{c},a_{t}^{c})-\alpha D(s^{c}_{t},s^{c}_{\text{target}})$ (15)
$=\mathbb{E}_{s_{t+1}^{c}}\left[r_{c}(s_{t}^{c})-\alpha\Delta(s_{t}^{c},s_{t+1}^{c})\right.$
$\left.+\underset{a_{t+1}^{c}\in\mathcal{A}}{\text{max}}(Q^{*}(s^{c}_{t+1},a^{c}_{t+1})-\alpha D(s^{c}_{t+1},s^{c}_{\text{target}}))\right].$
The $D(s^{c}_{t},s^{c}_{\text{target}})$ we define is only related to the state, so Eq.(15) is equivalent to Eq.(14). Notice that $r_{c}(s_{t}^{c})-\alpha\Delta(s_{t}^{c},s_{t+1}^{c})$ is the new reward function $\hat{r}_{c}(s^{c}_{t})$ we designed for GIANT. Hence, Eq.(15) can be viewed as
@@
$=\mathbb{E}_{s_{t+1}^{c}}\left[\hat{r}_{c}(s^{c}_{t})+\underset{a_{t+1}^{c}\in\mathcal{A}}{\text{max}}\hat{Q}^{*}(s^{c}_{t+1},a^{c}_{t+1})\right],$
where $\hat{Q}^{*}(s^{c}_{t},a^{c}_{t})=Q^{*}(s_{t}^{c},a_{t}^{c})-\alpha D(s^{c}_{t},s^{c}_{\text{target}})$. Eq.(16) is the Bellman optimal equation in dynamic path feedback. It constructs a new Markov Decision Process (MDP), which we use $M^{\prime}$ to denote for distinction and $M$ for the original MDP. Thus, the optimal policy in $M^{\prime}$ is
$\pi^{*}_{M^{\prime}}(s_{t}^{c})$ $=\text{arg}\underset{a_{t}^{c}\in\mathcal{A}}{\text{max}}\hat{Q}^{*}(s_{t}^{c},a_{t}^{c})$ (17)
$=\text{arg}\underset{a_{t}^{c}\in\mathcal{A}}{\text{max}}Q^{*}(s_{t}^{c},a_{t}^{c})-\alpha D(s^{c}_{t},s^{c}_{\text{target}})$
$=\text{arg}\underset{a_{t}^{c}\in\mathcal{A}}{\text{max}}Q^{*}(s_{t}^{c},a_{t}^{c})=\pi^{*}_{M}(s_{t}^{c}),$
which implies that GIANT learns optimal policy in dynamic path feedback is consistent with optimalcpolicy in the default rewards.

### v3

@@ B.2  Case Stuides
B.3  Cluster State Similarity
To assess the exploration efficiency of GIANT, we introduce cluster state similarity (CSS) as a metric, and in the main content we compare the performance of CURL and FULORA on only three tasks due to space limitations. Figure 7 also compare the model performance on remaining six tasks. For the same task, CURL experiences significant oscillations as path length increases, while FULORA remains stable. As in our previous analysis, CURL can only judge path quality by whether it has gone to the correct target cluster, while FULORA utilizes dynamic path feedback to promote the GIANT to converge to a high-quality path.
Figure 8: The ratios of the mean and variance of CSC difference between FULORA and CURL on different task-path_length.
Figure 8 further visualizes performances of FULORA and CURL on these tasks with varied path length. We utilize the ratio of the mean to the variance of CSS as an evaluation metric for the mean measures reasoning accuracy and the variance measures reasoning stability. In the vast majority of cases, FULORA outperforms CURL. In the case of long-distance reasoning for complex tasks like PersonBorninLocation-Path9 (PBL-9), AthletePlaysinLeague-Path7 (APL-9), AthletePlaysSport-Path9 (APS-9), and TeamPlaysinLeague (TPL), the performance gap between FULORA and CURL is significant.
B.4  Additional Ablation Studies
Here, we conduct a series of ablation experiments designed to answer the following three research questions:
RQ1: Can attention mechanism enhance DWARF’s reasoning ability?
RQ2: Can dynamic path feedback accelerate DWARF’S learning speed via improving GIANT’S reasoning ability?
RQ3: Can efficient guidance-exploration method make the guidance-exploration trade-off?
With a slight abuse of abbreviation, we use ATTN, DPF and GE to denote attention mechanism, dynamic path feedback and the efficient guide-exploration method respectively. To highlight the effect of each component, we set FULORA as the baseline. In Figure 9, FULORA-X represents pulling module X out of the FULORA framework. We conduct experiments on three different scales of KG to make a fair comparison. It is worth noting that in order to describe DWARF reasoning accuracy in the whole training process, Entity State Similarity (ESS, be similar to CSS we mentioned above) is used as an evaluation index.
B.4.1  Attention Mechanism
@@
B.4.2  Dynamic Path Feedback
@@
B.4.3  Efficient Guidance-Exploration
The core of FULORA is the efficient guidance-exploration method, which contributes to make the guidance-exploration trade-off. In the previous Hierarchical RL-based reasoning method like CURL, there is a high degree of coupling between the two agents. We treat ESS and CSS as two time series because they change with training progress. Next, we use Granger Causality test (Diks and Panchenko 2006) to examine the degree of coupling between DWARF and GIANT.
@@ B.4.3  Efficient Guidance-Exploration
The results of the Granger Causality Test for FULORA and FULORA-GE are shown in Table 7. The efficient guidance-exploration method reduces the coupling between GIANT and DWARF, thus alleviating the affect of policy shift caused by cluster mapping. In particular, on the WN18RR and FB15K-237, FULORA performs better than the version without efficient guidance-exploration method due to the greatly reduced coupling between DWARF and GIANT. Certain tasks in WN18RR (sparse KG) require the agent to have excellent long-distance reasoning ability. However, due to the distribution deviation, even if the GIANT reaches the correct target cluster, the guidance provided by it may not be suitable for DWARF (for example, the path length is too long to reach the correct target entity). The high coupling also seriously affects GIANT, resulting in poor model performance.
Figure 10: The effect of different cluster size used by GIANT. We present the link prediction (query answering) performance on three real-world KG datasets.
@@
C  Sensitive Test on Cluster Size
In the previous experiments, we examine the primary components of FULORA and analyze their impact on reasoning ability. Given that GIANT walks directly on the cluster-level KG derived from the entity-level KG, it is essential to separately evaluate the implications of cluster size.
Figure 10 shows the MRR score of FULORA under different cluster number $N$ during training on three real-world KG datasets. We observe that variations in cluster sizes significantly influence the reasoning performance of FULORA, particularly when reasoning over long distances on WN18RR (sparse KG) and FB15K-237 (dense KG). If the cluster size is too small (e.g., $N=500$), GIANT must traverse a greater number of clusters. In the absence of higher-level agent to provide guidance, the reasoning process approximates the long-distance reasoning of a single agent. Conversely, when the cluster size is too large (e.g., $N=50$), the guidance offered by GIANT becomes overly general. Even if the guidance is accurate, it may still prove ineffective, causing DWARF to similarly approximate the behavior of a single agent during long-distance reasoning.
Appendix C D  Time Series Analysis to Efficient Guidance-Exploration
Before Granger Causality test, we firstly use Augmented Dickey-Fuller (ADF) test (Cheung and Lai 1995) to judge the stationarity of the series. The results as shown in Table 8. Next, we make D-order difference for the non-stationary time series, and then test the stationarity of the difference series. As illustrated in Figure 11, the time series is stationary after 1-order difference. Therefore, we demonstrate that these series are integrated of order one. To avoid the pseudo-regression phenomenon, we also need to do cointegration analysis of these series (Dougherty 2011). Here, we employ Johansen test (Ho and Sørensen 1996) based on maximum likelihood estimation for cointegration analysis.
Figure 11: Entity state similarity curves and their 1-order difference on WN18RR and FB15K-237 from FULORA and FULORA-GE.
In the Johansen cointegration test (Ho and Sørensen 1996), the 2-cointegration term critical values of {10%, 5%, 1%} are {2.705, 3.841, 6.635} respectively. The critical value 3.841 corresponding to 5% was selected as the boundary for testing. Table 9 shows the results of Johansen cointegration test for FULORA and FULORA-GE. It shows that the seven non-stationary series avoid the pseudo-regression problem, thus the result of Granger causality test is reasonable (which is shown in Appendix B.4.3).
@@ Appendix C D  Time Series Analysis to Efficient Guidance-Exploration
DWARF(-GE) $\leftrightarrow$ GIANT(-GE) 0.037 74.582 0.001 4.133
@@
Appendix D E  Proof of Dynamic Path Feedback
A major concern with dynamic path feedback is whether the optimal policy GIANT learns is consistent with the default rewards. Here we provide the following proof of Theorem 1.
Proof 1
Recall Equation 9 where the next state $s^{c}_{t+1}$ is used, which means that it implies the action $a^{c}_{t}$. The next state $s^{c}_{t+1}$ is the output of the current state $s^{c}_{t}$ and the current action $a^{c}_{t}$ to the transition function. In fact, $J(\theta_{\pi^{c}})$ is a function of state $s^{c}_{t}$ and action $a^{c}_{t}$. The same goes for reward settings in the default environment. Therefore, we use the Q-function to simplify our notation,
@@ Appendix D E  Proof of Dynamic Path Feedback
$=\mathbb{E}_{s_{t+1}^{c}}\Big{[}r_{c}(s_{t}^{c})+\underset{a_{t+1}^{c}\in\mathcal{A}}{\rm{max}}Q^{*}(s^{c}_{t+1},a^{c}_{t+1})\Big{]},$
@@
$Q^{*}(s_{t}^{c},a_{t}^{c})-\alpha\rm{Sim}(s^{c}_{t},s^{c}_{\rm{target}})$ (13)
$=\mathbb{E}_{s_{t+1}^{c}}\Big{[}r_{c}(s_{t}^{c})-\alpha\Delta(s_{t}^{c},s_{t+1}^{c})$
$+\underset{a_{t+1}^{c}\in\mathcal{A}}{\rm{max}}(Q^{*}(s^{c}_{t+1},a^{c}_{t+1})-\alpha\rm{Sim}(s^{c}_{t+1},s^{c}_{\rm{target}}))\Big{]}.$
The $\rm{Sim}(s^{c}_{t},s^{c}_{\rm{target}})$ we define is only related to the state, so Equation 12 is equivalent to Equation 13. Notice that $r_{c}(s_{t}^{c})-\alpha\Delta(s_{t}^{c},s_{t+1}^{c})$ is the new reward function $\hat{r}_{c}(s^{c}_{t})$ we designed for GIANT. Hence, Equation 12 can be viewed as
@@
$=\mathbb{E}_{s_{t+1}^{c}}\Big{[}\hat{r}_{c}(s^{c}_{t})+\underset{a_{t+1}^{c}\in\mathcal{A}}{\rm{max}}\hat{Q}^{*}(s^{c}_{t+1},a^{c}_{t+1})\Big{]},$
where $\hat{Q}^{*}(s^{c}_{t},a^{c}_{t})=Q^{*}(s_{t}^{c},a_{t}^{c})-\alpha D(s^{c}_{t},s^{c}_{\rm{target}})$. Equation 14 is the Bellman optimal equation in dynamic path feedback. It constructs a new Markov Decision Process (MDP), which we use $M^{\prime}$ to denote for distinction and $M$ for the original MDP. Thus, the optimal policy in $M^{\prime}$ is
$\pi^{*}_{M^{\prime}}(s_{t}^{c})$ $=\rm{arg}\underset{a_{t}^{c}\in\mathcal{A}}{\rm{max}}\hat{Q}^{*}(s_{t}^{c},a_{t}^{c})$ (15)
$=\rm{arg}\underset{a_{t}^{c}\in\mathcal{A}}{\rm{max}}Q^{*}(s_{t}^{c},a_{t}^{c})-\alpha\rm{Sim}(s^{c}_{t},s^{c}_{\rm{target}})$
$=\rm{arg}\underset{a_{t}^{c}\in\mathcal{A}}{\rm{max}}Q^{*}(s_{t}^{c},a_{t}^{c})=\pi^{*}_{M}(s_{t}^{c}),$
[…]
Appendix E F  Pseudocode for FULORA
We show FULORA’s training process in one episode.
Algorithm 1 FULORA Training Algorithm (one episode)
0:    Entity-level KG $\mathcal{G}^{e}$ and cluster-level KG $\mathcal{G}^{c}$; Initial policy networks parameters $\theta_{\pi^{e}}$ and $\theta_{\pi^{c}}$; Initial Lagrange multiplier parameter $\theta_{\lambda}$ ;Source entity and cluster nodes $e_{s}$ and $c_{s}$; Entity-level query $r_{q}$; Target entity and cluster nodes $e_{o}$ and $c_{o}$; Maximum path length $T$
0:    parameters $\theta_{\pi^{e}}$, $\theta_{\pi^{c}}$, $\theta_{\lambda}$
1:  for $t=0,...,T-1$ do
2:     Set default cluster-level reward $r_{c}=1$ if $c_{t}=c_{o}$ otherwise $r_{c}=0$
3:     Set default entity-level reward $r_{e}=1$ if $e_{t}=e_{o}$ otherwise $r_{e}=0$
5:     Compute $J(\theta_{\pi^{c}})$, $J(\theta_{\pi^{e}})$, $J(\theta_{\lambda})$ based on Equation 7-9.
6:  end for
8:  return  $\theta_{\pi^{e}}$, $\theta_{\pi^{c}}$, $\theta_{\lambda}$
In addtion, a common concern is how clusters are formed, therefore, we provide pseudocode to demonstrate cluster formation, which is presented below.
Algorithm 2 Cluster Formation and Batch Processing
0:    Raw KG $\mathcal{G}$; Input data file, vocabularies (entity_vocab, relation_vocab, cluster_vocab), batch size batch_size, entity-to-cluster mapping entity_id_to_cluster_mapping
0:    Generated batches with cluster-level relations.
1:  Initialize vocabularies and mappings from the input data file
2:  Parse the input file to extract triples ($e_{1}$,$r$,$e_{2}$)
3:  Map each entity to its corresponding cluster
4:  Create cluster relations $c_{r}=c_{1}\_c_{2}$ for each triple
5:  Store the triples and update entity-level and cluster-level mappings
6:  For each batch, randomly sample triples from the stored data
7:  Extract entities $e_{1},e_{2}$, relations $r$ and their cluster mappings $c_{1},c_{2}$
8:  Yield entity-level and cluster-level batch data for training or testing
9:  return  Entity-level and cluster-level batch data.